OpenStudy (anonymous):

What will be the perimeter and the area of the rectangle below if it is enlarged using a scale factor of 4.5? A rectangle is shown. The length of the rectangle is labeled as 8 cm, and the width is labeled as 6 cm. Perimeter = 46 cm, area = 131.25 cm2 Perimeter = 126 cm, area = 972 cm2 Perimeter = 46 cm, area = 972 cm2 Perimeter = 126 cm, area = 131.25 cm2

OpenStudy (anonymous):

how did you get the answer? @Loser66

OpenStudy (loser66):

P =? A=?

OpenStudy (anonymous):

Did you delete your last comment, I didnt get to see it. @Loser66

OpenStudy (loser66):

Just answer me, Perimeter = ?? Area =?? (formula)

OpenStudy (anonymous):

Perimeter = 2(l+w) Area= l w is that what you are asking for? @Loser66

OpenStudy (loser66):

Good, now if the new l, namely l', l' = 4.5l and the new w , namely w' = 4.5w then the perimeter of the new one is 2(l' + w') , right? , replace the numbers you can get what you want. Same as area

OpenStudy (anonymous):

Ok, but how is the l and w the same? (4.5) I'm confused

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

They are not the same. just the ratio is the same. l = 8--> l' = 4.5 *8 w = 6, w' = 4.5*6